A 40.0 g sample of ethanol (Csp = 2.46 J/g °C) releases 2952 J as it cools from 50.0 °C. Calculate the change in temperature of the ethanol. _ °C
1
Expert's answer
2019-05-14T03:40:19-0400

Solution.

"Q = c \times m \times \Delta t = c \times m \times (t2-t1)"

"\Delta t = \frac{Q}{c \times m}"

"\Delta t = \frac{2952}{2.46\times40} = 30"

"\Delta t = t2-t1"

"t2 = \Delta t + t1"

t2 = 30 + 50 = 80

Answer:

t2 = 80